Fermi National Laboratory
 

Accelerator Update for May 23, 2007 - May 25, 2007

 

Wednesday May 23

The day shift began with Operations monitoring the Tevatron (TeV) store 5455 (68.01E30) and the Recycler stash (162.E10), with the Antiproton Source (Pbar) stacking (42.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI taking beam, with and MiniBooNE off for a scheduled access into MI-13, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

The MiniBooNE HVAC work began at 8:42 AM and ended at 9:45 AM. Operations began sending the experiment beam three minutes later.

Operations reported at 10:59 AM that power supply testing in the Meson enclosures had been completed.

Operations reported at 12:25 PM that interlock technicians completed the safety system testing for the Meson M02 M05 enclosures.

At 12:47 PM, a Booster vacuum pump (IPRF23) tripped off and caused two Booster RF stations (BRF13 & 14) to trip off. Everything reset.

A Pbar expert began a 20-minute stacktail study at 1:23 PM.

A MI expert began a multibatch slipstacking study at 2:03 PM.

Operations had the Meson M02 M05 enclosures ready for access at 3:20 PM.

A Pbar expert began a cooling study at 3:26 PM.

Operations reported at 4:09 PM that the Recycler had stashed antiprotons three times during the day shift.

The Pbar studies ended at 4:26 PM.

At 7:25 PM, the MI beam permit tripped off due to a bad beam position on the MI-8 collimators. It reset.

Operations terminated store 5455 at 8:59 PM.

Operations put the TeV into shot setup at 9:18 PM.

At 9:55 PM, Operations reported that they had to stop sending beam to MiniBooNE, NuMI, and Pbar while loading protons in the TeV due to bad beam positions in Booster.

At 10:08 PM, a Linac Klystron RF station (KRF5) tripped off. It reset.

At 10:17 PM, LRF3 tripped off. It reset.

Operations resumed sending beam to MiniBooNE, NuMI, and Pbar at 10:25 PM.

Operations established store 5456 at 11:24 PM with an initial luminosity of 240.47E30.

At 11:50 PM, MI tripped off due to the MI-8 collimators. It reset.

Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.

Thursday May 24

The midnight shift began with Operations monitoring store 5456 (212.13E30) and the stash (47.8E10), with Pbar stacking (16.2E10), and with NuMI and MiniBooNE operating normally.

At 1:02 AM, a Pbar Accumulator RF station (ARF1) tripped off for the first of two times in about three hours. It reset.

At 2:30 AM, LRF3 tripped off. It reset.

Interlock technicians began testing the Meson M01 safety system at 8 AM.

Operations reported at 8:05 AM that the Recycler had stashed three times during the midnight shift.

Operations reported at 8:27 AM that the Duty Electrician had turned off power for the Meson M02 enclosure.

At 9:13 AM, ARF1 tripped off. It reset.

Interlock techs reported at 9:17 AM that they had completed their M01 safety system test.

LRF3 tripped off at 12:27 PM. It reset.

Operations reported at 3:59 PM that the Recycler had stashed twice during the day shift.

Pbar experts reported at 4:46 PM that they had completed a reverse proton study.

At 5:02 PM, the Pbar lithium lens power supply tripped off. Operations put a NuMI only time line into the Time Line Generator (TLG) and called in a power supply expert.

Per Pbar expert instruction, Operations resumed stacking at 6:15 PM, but the lithium lens tripped off 10 minutes later. Operations resumed the NuMI only time line and called in more experts.

Operations prepared AP0 for an access at 8:16 PM so that experts could further investigate the lens problem.

At 11:19 PM, a Booster vacuum pump (IPRF22) tripped off and caused BRF8 to trip off.

Operations reported at 11:46 PM that mechanical support technicians were pulling out the lithium lens module and troubleshooting it. The lens module would be left out until morning when a second crew of mechanical techs would either repair or replace the module. Pbar can not stack without the lithium lens.

At midnight, Operations bypassed BRF8 and brought BRF7 online in its place.

Operations reported at midnight that the Recycler had stashed once during the evening shift.

Friday May 25

The midnight shift began with Operations monitoring store 5456 (45.4E30), the stash (243.8E10), and the stack (16.9E10), but with Pbar not stacking due to a problem with the lithium lens, and with NuMI and MiniBooNE taking beam.

At 12:50 AM, LRF3 tripped off for the first of two times in an hour. It reset.

Operations reported at 3:10 AM that the mechanical support crew completed pulling out the lithium lens. They also reported that the lens would have to be replaced due to a water leak on the lens transformer.

At 3:44 AM, LRF3 tripped off and would not reset. Operators contacted experts who then came in to investigate.

At 6:11 AM, a Linac expert reported that he found a short on the LRF3 driver grid. At 6:28 AM, he reported that the driver would have to be replaced.

The Plans for Friday and the Weekend

The plans for today are to replace the lithium lens and the LRF3 driver. During this work, the Run Coordinator will allow certain maintenance jobs as well as a study or two. The weekend will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/25/2007    email Fermilab